### v2

# LSEAttention is All You Need for Time Series Forecasting

Abstract
Transformer-based architectures have achieved remarkable success in natural language processing and computer vision. However, their performance in multivariate long-term forecasting often lags behind simpler linear baselines. Previous studies have identified the traditional attention mechanism as a significant factor contributing to this limitation. To unlock the full potential of transformers for multivariate time series forecasting, I introduce LSEAttention, an approach designed to address entropy collapse and training instability commonly observed in transformer models. I validate the effectiveness of LSEAttention across various real-world multivariate time series datasets, demonstrating that it not only outperforms existing time series transformer models but also exceeds the performance of some state-of-the-art models on specific datasets.
1 Introduction
Multivariate time series forecasting is crucial in diverse domains such as finance, healthcare, and environmental monitoring, where the goal is to predict future values based on historical data. This task is particularly challenging in long-term forecasting, as it necessitates models that can effectively capture feature correlations and long-term dependencies across multiple time series. Recent research has increasingly focused on applying transformer architectures to time series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms. However, despite their potential, state-of-the-art approaches for multivariate time series forecasting still primarily rely on linear models, such as those proposed by Ni et al. [1], raising concerns about the efficacy of transformers in this context, particularly given their suboptimal performance and inherent complexity.
1.1 Transformer-based Long-term Time Series Forecasting
The application of transformers to time series forecasting gained momentum following the introduction of the Transformer model [2] which established its potential as a universal backbone for various tasks across multiple modalities. The attention mechanism at the core of this architecture is formulated as follows:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
where $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively, and $d_{k}$ denotes the dimension of the key vectors. This formulation allows transformers to dynamically weigh the significance of various input elements, facilitating the capture of intricate dependencies within the data.
Recent studies have tailored transformer models specifically for long-term time series forecasting. For instance, Informer [3] introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys for forecasting, thereby reducing computational complexity and enhancing interpretability by focusing on essential temporal signals. Similarly, Autoformer [4] employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.
FEDformer [5] incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences. Instead of concentrating solely on point-wise attention, PatchTST [6] emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, which enables the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.
The challenges associated with applying transformers to time series forecasting are further highlighted by Ilbert et al. [7] in their SAMformer model. SAMformer integrates Reversible Instance Normalization (RevIN) [8] to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability. This approach tackles critical issues, including entropy collapse and training instability, that have impeded transformer models in this domain.
1.2 Entropy Collapse
In computer vision and natural language processing, attention matrices can experience entropy or rank collapse, as demonstrated by Dong et al. [9]. This issue is exacerbated in time series forecasting due to the frequent fluctuations inherent in temporal data, leading to substantial performance degradation. Ilbert et al. [7] identify the attention mechanism as a primary contributor to this problem and introduce SAM to mitigate it, thus enhancing model performance. However, the fundamental causes of entropy collapse remain inadequately addressed in the literature, warranting further exploration of its underlying mechanisms and effects on model performance.
1.3 Contributions
In this work, I present the following contributions:
  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the softmax operation.
  • Introduction of the LSEAttention module, a variation of traditional multi-head attention that addresses the entropy collapse issue, thereby enhancing the trainability and performance of transformers for time series forecasting.
  • Empirical validation of my approach on widely used multivariate long-term forecasting datasets, demonstrating that transformers equipped with LSEAttention outperform state-of-the-art models.
2 Proposed Approach
2.1 Problem Formulation
In multivariate time series forecasting, the objective is to predict future values $P$ for each channel, represented as $Y\in\mathbb{R}^{C\times P}$. This prediction is based on historical time series data of length $L$ across $C$ channels, encapsulated in the input matrix $X\in\mathbb{R}^{C\times L}$. The goal is to train a predictive model $f_{\omega}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times P}$, parameterized by $\omega$, to minimize the mean squared error (MSE) between the predicted and actual values. The MSE is defined as:
$\text{MSE}=\frac{1}{C}\sum_{c=1}^{C}\left\|Y_{c}-f_{\omega}(X_{c})\right\|^{2},$ (2)
where $Y_{c}$ and $X_{c}$ denote the true future values and historical input sequences for each channel $c$.
2.2 Motivation
Transformers depend heavily on their point-wise attention mechanism to capture temporal associations. However, this reliance can lead to a phenomenon known as attention collapse, where the attention matrices converge to similar values across inputs and prone to identical matrices, resulting in poor generalization. The underlying issue of entropy collapse remains underexplored, particularly concerning the usgae of transformer on time-series forecasting. Since the conditional number of softmax is defined similar to LSE (Log-Sum-Exp) which hints the numerical instability issue might be the core reason of attention collapse.
The condition number of a function quantifies its sensitivity to small input changes; a high condition number indicates that slight perturbations can lead to significant variations in output. For the softmax function, defined as:
$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}},$ (3)
where $x_{i}$ represents the $i$-th element of the input vector and $n$ is the total number of elements, the condition number is notably high. This instability can manifest as issues such as over-attention or entropy collapse, characterized by attention matrices with excessively high diagonal values (indicating overflow) and extremely low non-diagonal values (indicating underflow) .
To address these challenges, I propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) trick proposed by Blanchard et al. [10] along with the Gaussian Error Linear Unit (GELU) activation function introduced by Hendrycks et al. [11]. The LSE trick alleviates numerical instability arising from overflow and underflow through normalization. The softmax function can be reformulated using LSE as follows:
$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{e^{\text{LSE}(x)}},$ (4)
where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing numerical stability.
Normalizing the LSE yields:
$y=\log\left(e^{a}\prod_{i=1}^{n}e^{x_{i}-a}\right),$ (5)
where $a$ is a constant and max value utilized for normalization in practice. This simplifies to:
$y=a+\log\left(\prod_{i=1}^{n}e^{x_{i}-a}\right),$ (6)
and can be reformulated in a numerically stable manner as:
$g_{j}=\frac{e^{x_{j}-a}}{\sum_{i=1}^{n}e^{x_{i}-a}},\quad j=1,\ldots,n,$ (7)
where $g_{j}$ represents the $j$-th component of the stabilized softmax output.
This can be further expressed as:
$g_{j}=\exp\left(x_{j}-\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)\right),$ (8)
where $\sum_{i=1}^{n}e^{x_{i}}$ serves as the normalization factor, ensuring that the outputs sum to one.
Additionally, the GELU activation function provides smoother probabilistic activations, which help stabilize extreme log probability values before applying the exponential function, thereby mitigating abrupt transitions in attention scores. By approximating the ReLU function [12] with a smooth curve incorporating the cumulative distribution function (CDF) of the standard normal distribution, GELU reduces sudden shifts in activations that can occur with traditional ReLU. This property is particularly advantageous for stabilizing transformer-based attention mechanisms, where abrupt changes in activation can lead to numerical instability and gradient explosion.
The GELU function is formally defined as:
$\text{GELU}(x)=x\cdot\Phi(x),$ (9)
where $\Phi(x)$ represents the CDF of the standard normal distribution. This formulation ensures that GELU applies varying levels of scaling to inputs based on their magnitude, reducing the amplification of extreme values. The smooth, probabilistic nature of GELU allows for a gradual transition of input values, which, in turn, mitigates the impact of large gradient fluctuations during training.
This property becomes essential when combined with the Log-Sum-Exp (LSE) trick, which normalizes the softmax function in a numerically stable manner. Together, LSE and GELU prevent overflow and underflow in the exponential operations within softmax, resulting in a stabilized range of attention scores. This synergy enhances the robustness of transformer models by ensuring that attention weights remain well-distributed across tokens, ultimately leading to more stable gradients and improved convergence during training.
In traditional transformer architectures, the ReLU (Rectified Linear Unit) activation function used in the Feed-Forward Network (FFN) is susceptible to the ”dying ReLU” problem, where neurons can become inactive by outputting zero for all negative input values. This leads to a zero-gradient scenario for those neurons, effectively stalling their learning process and contributing to instability during training.
To address these challenges, Parametric ReLU (PReLU) is employed as an alternative activation function. PReLU introduces a learnable slope for negative inputs, allowing for a non-zero output even when the input is negative. This adaptation not only alleviates the dying ReLU problem but also facilitates a smoother transition between negative and positive activations, thereby enhancing the model’s ability to learn from all regions of the input space [13]. The presence of a non-zero gradient for negative values promotes improved gradient flow, which is crucial for training deeper architectures. Consequently, the use of PReLU contributes to overall training stability and helps maintain active representations, ultimately leading to enhanced model performance.
2.3 LATST: Overall Structure
The proposed LSEAttention Time-Series Transformer (LATST) Fig. 1 builds upon the modifications discussed in Motivation Sec. 2, introducing a key enhancement: Reversible Instance Normalization. This normalization technique is particularly effective for mitigating discrepancies between training and testing data distributions in time series forecasting tasks as in [7]
The architecture retains a traditional temporal self-attention mechanism, integral to the LSEAttention module. To assess the efficacy of the LSEAttention component, I conducted an ablation study comparing LATST with the Gaussian Error Linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified Linear Unit (ReLU). The details of this study can be found in Ablation Study of Sec. 3.
Overall, the LATST architecture consists of a single-layer transformer framework augmented with substitution modules, allowing for adaptive learning while maintaining the robustness of attention mechanisms. This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
Figure 1: Overall structure of LATST
3 Experiments
3.1 Dataset
I conduct experiments on eight publicly available datasets of real-world multivariate time series that are widely recognized for long-term forecasting. These datasets include four Electricity Transformer Temperature datasets: ETTh1, ETTh2, ETTm1, and ETTm2, as well as the Electricity, Weather and Traffic datasets are sourced from Wu et al. [4]. The Exchange dataset is from Lai et al. [14]. Each time series is segmented based on a specified input length to facilitate effective modeling.
Sequence and Prediction Length: In accordance with prior works, I set the input sequence length at 336 time steps. The prediction length, which indicates the number of future time steps for which I aim to forecast data, varies among the set {96, 192, 336, 720}. This flexibility in prediction length allows for an evaluation of the model’s performance across different forecasting horizons.
Evaluation Metric: Consistent with previous studies, I employ the widely used evaluation metric, mean squared error (MSE). This metric has been a standard choice in earlier works and is defined as follows:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(Y_{i}-\hat{Y}_{i})^{2},$ (10)
where $N$ is the total number of samples, $Y_{i}$ is the $i$-th prediction, and $\hat{Y}_{i}$ represents the corresponding ground truth value.
3.2 Baseline
I compare the performance of LATST against leading transformer-based models, including SAMformer [7], TSMixer [15], iTransformer [16], and PatchTST [6].
Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
Hyperparameter Values
Batch size 8
Initial learning rate 0.5 to 0.00001
Number of heads 4 to 16
Head dropout rate 0, 0.2, 0.3
Input sequence length 336
Prediction length 96, 192, 336, 720
Table 1: Hyperparameter settings used in experiments.
Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons. The best results are highlighted in bold. LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while results for other models are sourced from Ilbert et al. [7].
Methods LATST SAMformer Transformer TSMixer iTrans PatchTST
Electricity 96 0.131 0.155 0.182 0.173 - -
192 0.154 0.168 0.202 0.204 - -
336 0.169 0.183 0.212 0.217 - -
720 0.197 0.219 0.238 0.242 - -
Traffic 96 0.380 0.407 0.420 0.409 0.395 0.462
192 0.406 0.415 0.441 0.637 0.417 0.466
336 0.417 0.421 0.501 0.747 0.433 0.482
720 0.455 0.456 0.468 0.688 0.467 0.514
Weather 96 0.156 0.197 0.227 0.214 0.174 0.177
192 0.202 0.235 0.256 0.231 0.221 0.225
336 0.254 0.276 0.278 0.279 0.278 0.278
720 0.329 0.334 0.353 0.343 0.358 0.354
ETTh1 96 0.390 0.381 0.509 0.398 0.386 0.414
192 0.426 0.409 0.535 0.426 0.441 0.460
336 0.414 0.423 0.570 0.435 0.487 0.501
720 0.426 0.427 0.601 0.498 0.503 0.500
ETTh2 96 0.290 0.295 0.396 0.308 0.297 0.302
192 0.350 0.340 0.413 0.352 0.380 0.388
336 0.380 0.350 0.414 0.360 0.428 0.426
720 0.399 0.391 0.424 0.409 0.427 0.431
ETTm1 96 0.300 0.329 0.384 0.336 0.334 0.329
192 0.345 0.353 0.400 0.362 0.377 0.367
336 0.375 0.382 0.461 0.391 0.426 0.399
720 0.431 0.429 0.463 0.450 0.491 0.454
ETTm2 96 0.176 0.181 0.200 0.211 0.180 0.175
192 0.231 0.233 0.273 0.252 0.250 0.241
336 0.280 0.285 0.310 0.303 0.311 0.305
720 0.376 0.375 0.426 0.390 0.412 0.402
Exchange 96 0.098 0.161 0.292 0.343 0.086 0.088
192 0.242 0.246 0.372 0.342 0.177 0.176
336 0.408 0.368 0.494 0.484 0.331 0.301
720 0.898 1.003 1.323 1.204 0.847 0.901
3.3 Performance Analysis
LATST demonstrates superior performance by achieving the best mean squared error (MSE) scores in 21 out of 32 scenarios across the evaluated eight datasets. This remarkable consistency underscores the model’s robustness and adaptability in addressing diverse time series forecasting tasks. Specifically, LATST outperforms all other transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, indicating its efficacy in managing complex temporal dependencies.
When comparing LATST directly with SAMformer, LATST exhibits an overall improvement of 4.3% in MSE. Furthermore, LATST surpasses the Transformer model by a substantial margin of 19.11%, highlighting its enhanced predictive capabilities. The modifications introduced in LATST, including the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute to this performance enhancement, allowing the model to better capture long-range dependencies and temporal patterns.
The detailed performance metrics for each dataset and prediction length are summarized in Tab. 2. Notably, LATST consistently achieves the lowest MSE across multiple input lengths for datasets such as ETTh1, ETTh2, and Electricity. For instance, in the Electricity dataset, LATST records an MSE of 0.131 for a prediction length of 96, significantly outperforming the closest competitor, SAMformer, which achieves an MSE of 0.155. This trend is replicated across various prediction horizons, demonstrating the model’s ability to maintain stability and robustness under different forecasting conditions.
Consistency Across Datasets: The model’s ability to achieve the best MSE scores in a significant number of scenarios reinforces its robustness and generalizability to diverse forecasting challenges. The performance stability observed across different datasets suggests that LATST effectively learns and captures underlying temporal patterns, enabling it to adapt well to varying data distributions and dynamics. For example, in the Traffic dataset, LATST maintains lower MSE values even at longer prediction horizons, indicating its resilience against overfitting and its capacity to generalize beyond training conditions.
3.4 Ablation Study
In this section, I present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework. By systematically removing or modifying key elements, such as the GeLU and the PReLU, I isolate their contributions to the overall performance of the model. For consistency, I fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
Dataset Pred. Length With PReLU With ReLU
ETTh2 96 0.290 0.292
ETTm1 96 0.300 0.305
ETTm2 96 0.176 0.191
Exchange Rate 96 0.098 0.115
Table 3: Performance of LATST with different activation functions
Preliminary results from the ablation study, as shown in Tab. 3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets. This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
To further investigate the contribution of the Log-Sum-Exp trick, I conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
Dataset Pred. Length Without GeLU Transformer
ETTh2 96 0.312 0.396
ETTm1 96 0.333 0.384
ETTm2 96 0.183 0.200
Exchange Rate 96 0.105 0.292
Table 4: Performance of LATST without GeLU compared to Transformer
As illustrated in Tab. 4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration. This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function. The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
These findings not only elucidate the effectiveness of various components within the LATST architecture but also provide valuable insights for future enhancements. I aim to explore optimal configurations of these modules to further refine the model’s performance and expand its applicability to more complex time series forecasting challenges.
4 Conclusion and Future Work
In this paper, I present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting. Our empirical results confirm the effectiveness and robustness of the proposed LATST approach across a variety of forecasting contexts. Although I do not exhaustively explore all potential causes of entropy collapse, I introduce a novel module specifically designed to address this issue.
Looking ahead, future research will focus on a comprehensive investigation of the remaining factors contributing to entropy collapse. This exploration aims to identify additional modifications and enhancements that could further improve the stability and performance of Transformer models in time series forecasting.
In summary, I believe that our work not only addresses critical challenges associated with the entropy collapse phenomenon but also sets the stage for future advancements in time series forecasting methodologies.

### v3

@@
 Transformer-based architectures have achieved remarkable success in natural language processing and computer vision.
-However, their performance in multivariate long-term forecasting often lags behind simpler linear baselines.
-Previous studies have identified the traditional attention mechanism as a significant factor contributing to this limitation.
-To unlock the full potential of transformers for multivariate time series forecasting, I introduce LSEAttention, an approach designed to address entropy collapse and training instability commonly observed in transformer models.
-I validate the effectiveness of LSEAttention across various real-world multivariate time series datasets, demonstrating that it not only outperforms existing time series transformer models but also exceeds the performance of some state-of-the-art models on specific datasets.
+However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines.
+Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting.
+We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models.
+Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
 1 Introduction
@@
 This task is particularly challenging in long-term forecasting, as it necessitates models that can effectively capture feature correlations and long-term dependencies across multiple time series.
-Recent research has increasingly focused on applying transformer architectures to time series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms.
-However, despite their potential, state-of-the-art approaches for multivariate time series forecasting still primarily rely on linear models, such as those proposed by Ni et al.
+Recent research has increasingly focused on applying Transformer architectures to time-series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms.
+However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Zenget al.
 [1], raising concerns about the efficacy of Transformers in this context, particularly given their suboptimal performance and inherent complexity.
 1.1 Transformer-based Long-term Time Series Forecasting
-The application of transformers to time series forecasting gained momentum following the introduction of the Transformer model [2] which established its potential as a universal backbone for various tasks across multiple modalities.
-The attention mechanism at the core of this architecture is formulated as follows:
+The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model which established its potential as a universal backbone for various tasks across multiple modalities.[2] The attention mechanism at the core of this architecture is formulated as follows:
 $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
@@
 This formulation allows Transformers to dynamically weigh the significance of various input elements, facilitating the capture of intricate dependencies within the data.
-Recent studies have tailored transformer models specifically for long-term time series forecasting.
-For instance, Informer [3] introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys for forecasting, thereby reducing computational complexity and enhancing interpretability by focusing on essential temporal signals.
-Similarly, Autoformer [4] employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.
-FEDformer [5] incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.
-Instead of concentrating solely on point-wise attention, PatchTST [6] emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, which enables the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.
-The challenges associated with applying transformers to time series forecasting are further highlighted by Ilbert et al.
-[7] in their SAMformer model.
-SAMformer integrates Reversible Instance Normalization (RevIN) [8] to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.
-This approach tackles critical issues, including entropy collapse and training instability, that have impeded transformer models in this domain.
+Recent studies have tailored Transformer models specifically for long-term time-series forecasting.
+For example, Informer introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.[3] Similarly, Autoformer employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.[4]
+FEDformer incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.[5] Instead of concentrating solely on point-wise attention, PatchTST emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.[6]
+The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert in their SAMformer model.
+SAMformer integrates Reversible Instance Normalization (RevIN) to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.et al.
+[8, 7]This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
 1.2 Entropy Collapse
@@
 [9].
-This issue is exacerbated in time series forecasting due to the frequent fluctuations inherent in temporal data, leading to substantial performance degradation.
+This issue is exacerbated in time-series forecasting due to the frequent fluctuations inherent in temporal data, leading to substantial performance degradation.
 Ilbert et al.
-[7] identify the attention mechanism as a primary contributor to this problem and introduce SAM to mitigate it, thus enhancing model performance.
+[7] identified the attention mechanism as a primary contributor to this problem and introduced SAM to mitigate it, thus enhancing model performance.
 However, the fundamental causes of entropy collapse remain inadequately addressed in the literature, warranting further exploration of its underlying mechanisms and effects on model performance.
 1.3 Contributions
-In this work, I present the following contributions:
-  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the softmax operation.
-  • Introduction of the LSEAttention module, a variation of traditional multi-head attention that addresses the entropy collapse issue, thereby enhancing the trainability and performance of transformers for time series forecasting.
-  • Empirical validation of my approach on widely used multivariate long-term forecasting datasets, demonstrating that transformers equipped with LSEAttention outperform state-of-the-art models.
+In this work, we present the following contributions:
+  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the attention for time-series data.
+  • Introduction of the LATST (LESAttention Time-Series Transformer) a Transformer with LSEAttention replacement that attempts to solve the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
+  • Empirical validation of our approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art Transformers.
 2 Proposed Approach
@@
 2.2 Motivation
-Transformers depend heavily on their point-wise attention mechanism to capture temporal associations.
-However, this reliance can lead to a phenomenon known as attention collapse, where the attention matrices converge to similar values across inputs and prone to identical matrices, resulting in poor generalization.
-The underlying issue of entropy collapse remains underexplored, particularly concerning the usgae of transformer on time-series forecasting.
-Since the conditional number of softmax is defined similar to LSE (Log-Sum-Exp) which hints the numerical instability issue might be the core reason of attention collapse.
-The condition number of a function quantifies its sensitivity to small input changes; a high condition number indicates that slight perturbations can lead to significant variations in output.
-For the softmax function, defined as:
-$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}},$ (3)
-where $x_{i}$ represents the $i$-th element of the input vector and $n$ is the total number of elements, the condition number is notably high.
-This instability can manifest as issues such as over-attention or entropy collapse, characterized by attention matrices with excessively high diagonal values (indicating overflow) and extremely low non-diagonal values (indicating underflow) .
-To address these challenges, I propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) trick proposed by Blanchard et al.
+Transformers have emerged as a powerful architecture for capturing temporal associations in sequential data, primarily due to their reliance on point-wise attention mechanisms.
+However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly identical matrix.
+This identity diminishes the model’s ability to differentiate patterns, thereby hindering generalization.
+An equally critical yet less-explored issue is entropy collapse, particularly prevalent in the application of Transformers to time-series forecasting.
+In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation.[1] A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
+In the attention mechanism Eq.
+1, attention scores are computed as the dot product of query and key vectors.
+This operation tends to disproportionately emphasize abrupt changes or noise by assigning them larger values.
+Subsequently, when these scores are passed through the softmax function, defined as:
+$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{i}}},$ (3)
+The exponential computation introduces significant numerical instability.
+Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.[10] This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation.
+Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output.
+This disproportionate weighting undermines the diversity of the attention mechanism, reducing its capacity to effectively model complex temporal relationships.
+As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data which is the inherent characteristic of time-series data.
+This behavior culminates in entropy collapse, where the attention distribution becomes overly concentrated on a few elements, thereby limiting the model’s representational power.
+In other words, trend and periodicity might be barely considered.[1] Addressing this issue necessitates exploring alternatives to the conventional softmax function.
+$\text{LSE}(x)=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right).$ (4)
+To address these challenges, we propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eq.
+4 trick proposed by Blanchard et al.
 [10] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al.
-[11].
-The LSE trick alleviates numerical instability arising from overflow and underflow through normalization.
-The softmax function can be reformulated using LSE as follows:
-$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{e^{\text{LSE}(x)}},$ (4)
-where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing numerical stability.
-Normalizing the LSE yields:
-$y=\log\left(e^{a}\prod_{i=1}^{n}e^{x_{i}-a}\right),$ (5)
-where $a$ is a constant and max value utilized for normalization in practice. This simplifies to:
-$y=a+\log\left(\prod_{i=1}^{n}e^{x_{i}-a}\right),$ (6)
-and can be reformulated in a numerically stable manner as:
-$g_{j}=\frac{e^{x_{j}-a}}{\sum_{i=1}^{n}e^{x_{i}-a}},\quad j=1,\ldots,n,$ (7)
-where $g_{j}$ represents the $j$-th component of the stabilized softmax output.
-This can be further expressed as:
-$g_{j}=\exp\left(x_{j}-\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)\right),$ (8)
-where $\sum_{i=1}^{n}e^{x_{i}}$ serves as the normalization factor, ensuring that the outputs sum to one.
-Additionally, the GELU activation function provides smoother probabilistic activations, which help stabilize extreme log probability values before applying the exponential function, thereby mitigating abrupt transitions in attention scores.
-By approximating the ReLU function [12] with a smooth curve incorporating the cumulative distribution function (CDF) of the standard normal distribution, GELU reduces sudden shifts in activations that can occur with traditional ReLU.
-This property is particularly advantageous for stabilizing transformer-based attention mechanisms, where abrupt changes in activation can lead to numerical instability and gradient explosion.
-The GELU function is formally defined as:
-$\text{GELU}(x)=x\cdot\Phi(x),$ (9)
-where $\Phi(x)$ represents the CDF of the standard normal distribution.
-This formulation ensures that GELU applies varying levels of scaling to inputs based on their magnitude, reducing the amplification of extreme values.
-The smooth, probabilistic nature of GELU allows for a gradual transition of input values, which, in turn, mitigates the impact of large gradient fluctuations during training.
-This property becomes essential when combined with the Log-Sum-Exp (LSE) trick, which normalizes the softmax function in a numerically stable manner.
-Together, LSE and GELU prevent overflow and underflow in the exponential operations within softmax, resulting in a stabilized range of attention scores.
-This synergy enhances the robustness of transformer models by ensuring that attention weights remain well-distributed across tokens, ultimately leading to more stable gradients and improved convergence during training.
+[11] to form a gated attention .
+The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value.
+Starting from LSE where the softmax can be represented with LSE in this way:
+$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
+where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing the numerical stability.
+The LSE trick is stated as:
+$y=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)=\log\left(a\sum_{i=1}^{n}e^{x_{i}}-a\right)$ (6)
+where $a$ is a constant and maximum value utilized ensuring that the largest exponent becomes zero and the others become non-positive.
+As a result, all exponential terms are bounded between 0 and 1 like normalization for numerical stabilization purpose in practice.
+This can be simplified to:
+$y=a+\log\left(\sum_{i=1}^{n}e^{x_{i}-a}\right)$ (7)
+To enhance the model’s capabitlity, the GELU is applied to introduce the non-linearity, then the equation is modfied as:
+$y=\text{GELU}\left(a+\log\left(\sum_{i=1}^{n}e^{x_{i}-a}\right)\right)$ (8)
+This eventually can be expressed as the numerically stable manner of softmax function:
+$g_{i}=\exp(x_{i}-y)$ (9)
+where $g_{i}$ represents the $i$-th component of the stabilized softmax output.
 In traditional Transformer architectures, the ReLU (Rectified linear Unit) activation function used in the Feed-Forward Network (FFN) is susceptible to the ”dying ReLU” problem, where neurons can become inactive by outputting zero for all negative input values.
 This leads to a zero-gradient scenario for those neurons, effectively stalling their learning process and contributing to instability during training.
-To address these challenges, Parametric ReLU (PReLU) is employed as an alternative activation function.
+To address these challenges, Parametric ReLU (PReLU) is employed as an alternative activation function for Feed-Forward Network.
 PReLU introduces a learnable slope for negative inputs, allowing for a non-zero output even when the input is negative.
@@
 This normalization technique is particularly effective for mitigating discrepancies between training and testing data distributions in time series forecasting tasks as in [7]
-The architecture retains a traditional temporal self-attention mechanism, integral to the LSEAttention module.
-To assess the efficacy of the LSEAttention component, I conducted an ablation study comparing LATST with the Gaussian Error Linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified Linear Unit (ReLU).
+The architecture replaces the traditional temporal self-attention mechanism, with the LSEAttention module.
+To assess the efficacy of the LSEAttention component, we conducted an ablation study comparing LATST with the Gaussian Error linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified linear Unit (ReLU).
 The details of this study can be found in Ablation Study of Sec.
 3.
-Overall, the LATST architecture consists of a single-layer transformer framework augmented with substitution modules, allowing for adaptive learning while maintaining the robustness of attention mechanisms.
+Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing adaptive learning while maintaining the robustness of attention mechanisms.
 This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
-Figure 1: Overall structure of LATST
+Figure 1: Overall structure of LATST.
 3 Experiments
 3.1 Dataset
-I conduct experiments on eight publicly available datasets of real-world multivariate time series that are widely recognized for long-term forecasting.
+We conduct experiments on eight publicly available datasets of real-world multivariate time series that are widely recognized for long-term forecasting.
 These datasets include four Electricity Transformer Temperature datasets: ETTh1, ETTh2, ETTm1, and ETTm2, as well as the Electricity, Weather and Traffic datasets are sourced from Wu et al.
@@
 Each time series is segmented based on a specified input length to facilitate effective modeling.
-Sequence and Prediction Length: In accordance with prior works, I set the input sequence length at 336 time steps.
-The prediction length, which indicates the number of future time steps for which I aim to forecast data, varies among the set {96, 192, 336, 720}.
+Sequence and Prediction Length: In accordance with prior works, We set the input sequence length at 336 time steps.
+The prediction length, which indicates the number of future time steps for which We aim to forecast data, varies among the set {96, 192, 336, 720}.
 This flexibility in prediction length allows for an evaluation of the model’s performance across different forecasting horizons.
-Evaluation Metric: Consistent with previous studies, I employ the widely used evaluation metric, mean squared error (MSE).
+Evaluation Metric: Consistent with previous studies, we employ the widely used evaluation metric, mean squared error (MSE).
 This metric has been a standard choice in earlier works and is defined as follows:
@@
 3.2 Baseline
-I compare the performance of LATST against leading transformer-based models, including SAMformer [7], TSMixer [15], iTransformer [16], and PatchTST [6].
+We compare the performance of LATST against leading Transformer-based models, including SAMformer [7], TSMixer [15], iTransformer [16], and PatchTST [6].
 Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
+Table 1: Hyperparameter settings used in experiments.
 Hyperparameter Values
 Batch size 8
-Initial learning rate 0.5 to 0.00001
+Initial learning rate 0.5 to 0.00005
 Number of heads 4 to 16
@@
 Prediction length 96, 192, 336, 720
-Table 1: Hyperparameter settings used in experiments.
 Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
@@
 [7].
-Methods LATST SAMformer Transformer TSMixer iTrans PatchTST
-Electricity 96 0.131 0.155 0.182 0.173 - -
-192 0.154 0.168 0.202 0.204 - -
-336 0.169 0.183 0.212 0.217 - -
-720 0.197 0.219 0.238 0.242 - -
-Traffic 96 0.380 0.407 0.420 0.409 0.395 0.462
-192 0.406 0.415 0.441 0.637 0.417 0.466
-336 0.417 0.421 0.501 0.747 0.433 0.482
-720 0.455 0.456 0.468 0.688 0.467 0.514
-Weather 96 0.156 0.197 0.227 0.214 0.174 0.177
-192 0.202 0.235 0.256 0.231 0.221 0.225
-336 0.254 0.276 0.278 0.279 0.278 0.278
-720 0.329 0.334 0.353 0.343 0.358 0.354
-ETTh1 96 0.390 0.381 0.509 0.398 0.386 0.414
-192 0.426 0.409 0.535 0.426 0.441 0.460
-336 0.414 0.423 0.570 0.435 0.487 0.501
-720 0.426 0.427 0.601 0.498 0.503 0.500
-ETTh2 96 0.290 0.295 0.396 0.308 0.297 0.302
-192 0.350 0.340 0.413 0.352 0.380 0.388
-336 0.380 0.350 0.414 0.360 0.428 0.426
-720 0.399 0.391 0.424 0.409 0.427 0.431
-ETTm1 96 0.300 0.329 0.384 0.336 0.334 0.329
-192 0.345 0.353 0.400 0.362 0.377 0.367
-336 0.375 0.382 0.461 0.391 0.426 0.399
-720 0.431 0.429 0.463 0.450 0.491 0.454
-ETTm2 96 0.176 0.181 0.200 0.211 0.180 0.175
-192 0.231 0.233 0.273 0.252 0.250 0.241
-336 0.280 0.285 0.310 0.303 0.311 0.305
-720 0.376 0.375 0.426 0.390 0.412 0.402
-Exchange 96 0.098 0.161 0.292 0.343 0.086 0.088
-192 0.242 0.246 0.372 0.342 0.177 0.176
-336 0.408 0.368 0.494 0.484 0.331 0.301
-720 0.898 1.003 1.323 1.204 0.847 0.901
+Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
+ECL 96 0.131 0.155 0.182 0.173 - 0.140
+192 0.154 0.168 0.202 0.204 - 0.153
+336 0.169 0.183 0.212 0.217 - 0.169
+720 0.197 0.219 0.238 0.242 - 0.203
+Traffic 96 0.380 0.407 0.420 0.409 0.395 0.410
+192 0.406 0.415 0.441 0.637 0.417 0.423
+336 0.417 0.421 0.501 0.747 0.433 0.436
+720 0.455 0.456 0.468 0.688 0.467 0.466
+Weather 96 0.152 0.197 0.227 0.214 0.174 0.176
+192 0.196 0.235 0.256 0.231 0.221 0.220
+336 0.249 0.276 0.278 0.279 0.278 0.265
+720 0.317 0.334 0.353 0.343 0.358 0.323
+ETTh1 96 0.390 0.381 0.509 0.398 0.386 0.375
+192 0.426 0.409 0.535 0.426 0.441 0.405
+336 0.414 0.423 0.570 0.435 0.487 0.439
+720 0.426 0.427 0.601 0.498 0.503 0.472
+ETTh2 96 0.290 0.295 0.396 0.308 0.297 0.289
+192 0.348 0.340 0.413 0.352 0.380 0.383
+336 0.380 0.350 0.414 0.360 0.428 0.448
+720 0.399 0.391 0.424 0.409 0.427 0.605
+ETTm1 96 0.295 0.329 0.384 0.336 0.334 0.299
+192 0.345 0.353 0.400 0.362 0.377 0.335
+336 0.375 0.382 0.461 0.391 0.426 0.369
+720 0.431 0.429 0.463 0.450 0.491 0.425
+ETTm2 96 0.176 0.181 0.200 0.211 0.180 0.167
+192 0.229 0.233 0.273 0.252 0.250 0.224
+336 0.280 0.285 0.310 0.303 0.311 0.281
+720 0.376 0.375 0.426 0.390 0.412 0.397
+Exchange 96 0.098 0.161 0.292 0.343 0.086 0.081
+192 0.218 0.246 0.372 0.342 0.177 0.157
+336 0.408 0.368 0.494 0.484 0.331 0.305
+720 0.865 1.003 1.323 1.204 0.847 0.603
+Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
+Methods LATST SAMformer
+Prediction Length 96 192 336 720 96 192 336 720
+Electricity 251K 263K 282K 331K 82K 131K 205K 401K
+Traffic 251K 263K 282K 331K 82K 131K 205K 401K
+Weather 51K 54K 58K 72K 82K 131K 205K 401K
+ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
+ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
+ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
+ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
+Exchange 51K 115K 199K 236K 82K 131K 205K 401K
 3.3 Performance Analysis
-LATST demonstrates superior performance by achieving the best mean squared error (MSE) scores in 21 out of 32 scenarios across the evaluated eight datasets.
-This remarkable consistency underscores the model’s robustness and adaptability in addressing diverse time series forecasting tasks.
-Specifically, LATST outperforms all other transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, indicating its efficacy in managing complex temporal dependencies.
-When comparing LATST directly with SAMformer, LATST exhibits an overall improvement of 4.3% in MSE.
-Furthermore, LATST surpasses the Transformer model by a substantial margin of 19.11%, highlighting its enhanced predictive capabilities.
-The modifications introduced in LATST, including the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute to this performance enhancement, allowing the model to better capture long-range dependencies and temporal patterns.
+LATST demonstrates exceptional performance, achieving the lowest mean squared error (MSE) scores in 21 out of 32 scenarios across eight diverse datasets.
+This remarkable consistency underscores the model’s robustness and adaptability, effectively addressing a wide range of time-series forecasting tasks.
+Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, demonstrating its superior ability to manage complex temporal dependencies.
+When comparing LATST directly with SAMformer, the model shows its superiority by outperforming SAMformer on 6 out of 8 datasets with overall improvement by 5.12% .
+Moreover, LATST surpasses the traditional Transformer model by a notable margin of 19.75%, which emphasizes its enhanced predictive capabilities.
+Key innovations in LATST, such as the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute significantly to this performance enhancement.
+These modifications enable LATST to capture long-range dependencies and temporal patterns more effectively, setting it apart from its counterparts.
 The detailed performance metrics for each dataset and prediction length are summarized in Tab.
 2.
-Notably, LATST consistently achieves the lowest MSE across multiple input lengths for datasets such as ETTh1, ETTh2, and Electricity.
-For instance, in the Electricity dataset, LATST records an MSE of 0.131 for a prediction length of 96, significantly outperforming the closest competitor, SAMformer, which achieves an MSE of 0.155.
-This trend is replicated across various prediction horizons, demonstrating the model’s ability to maintain stability and robustness under different forecasting conditions.
-Consistency Across Datasets: The model’s ability to achieve the best MSE scores in a significant number of scenarios reinforces its robustness and generalizability to diverse forecasting challenges.
-The performance stability observed across different datasets suggests that LATST effectively learns and captures underlying temporal patterns, enabling it to adapt well to varying data distributions and dynamics.
-For example, in the Traffic dataset, LATST maintains lower MSE values even at longer prediction horizons, indicating its resilience against overfitting and its capacity to generalize beyond training conditions.
+Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variants.
+This phenomenon is closely related to the number of variants and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions.
+A more in-depth analysis can be found in Tab.
+3.
+In comparison with DLinear, LATST outperforms the model on 5 out of 8 datasets, despite DLinear having 139.7K parameters.
+LATST, with fewer parameters in most scenarios, demonstrates that its performance is not solely dependent on model size, but rather on the efficiency and effectiveness of its architecture.
+Consistency Across Datasets: LATST’s ability to achieve the best MSE scores in a significant number of scenarios further reinforces its robustness and generalizability across diverse forecasting challenges.
+This performance stability across different datasets suggests that LATST effectively learns and captures the underlying temporal patterns, allowing it to adapt well to varying data distributions and dynamics.
+For instance, in the Traffic dataset, LATST consistently maintains lower MSE values even at longer prediction horizons, illustrating its resilience against overfitting and its capacity to generalize beyond training conditions.
+Table 4: Performance of LATST with different activation functions for prediction length of 96.
+Dataset With PReLU With ReLU
+ETTh2 0.290 0.292
+ETTm1 0.295 0.305
+ETTm2 0.176 0.191
+Exchange Rate 0.098 0.115
 3.4 Ablation Study
-In this section, I present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework.
-By systematically removing or modifying key elements, such as the GeLU and the PReLU, I isolate their contributions to the overall performance of the model.
-For consistency, I fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
-Dataset Pred. Length With PReLU With ReLU
-ETTh2 96 0.290 0.292
-ETTm1 96 0.300 0.305
-ETTm2 96 0.176 0.191
-Exchange Rate 96 0.098 0.115
-Table 3: Performance of LATST with different activation functions
-Preliminary results from the ablation study, as shown in Tab.
-3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
+In this section, we present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework.
+By systematically removing or modifying key elements, such as the GeLU and the PReLU, their contributions are isolated to the overall performance of the model.
+For consistency, we fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
+Table 5: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
+Dataset Without GeLU Transformer
+ETTh2 0.312 0.396
+ETTm1 0.333 0.384
+ETTm2 0.183 0.200
+Exchange Rate 0.105 0.292
+Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
 This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
-To further investigate the contribution of the Log-Sum-Exp trick, I conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
-Dataset Pred. Length Without GeLU Transformer
-ETTh2 96 0.312 0.396
-ETTm1 96 0.333 0.384
-ETTm2 96 0.183 0.200
-Exchange Rate 96 0.105 0.292
-Table 4: Performance of LATST without GeLU compared to Transformer
-As illustrated in Tab.
-4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
+To further investigate the contribution of the Log-Sum-Exp trick, we conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
+As illustrated in LABEL:tab:ab_gelu, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
 This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function.
 The improvement achieved through the Log-Sum-Exp trick is attributed to its ability to stabilize gradient computations during training, thereby facilitating the learning of more complex patterns.
-These findings not only elucidate the effectiveness of various components within the LATST architecture but also provide valuable insights for future enhancements.
-I aim to explore optimal configurations of these modules to further refine the model’s performance and expand its applicability to more complex time series forecasting challenges.
+These findings not only elucidate the effectiveness of various components within the LATST architecture, but also provide valuable insights for future enhancements.
+We aim to explore optimal configurations of these modules to further refine the model’s performance and expand its applicability to more complex time series forecasting challenges.
 4 Conclusion and Future Work
-In this paper, I present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting.
-Our empirical results confirm the effectiveness and robustness of the proposed LATST approach across a variety of forecasting contexts.
-Although I do not exhaustively explore all potential causes of entropy collapse, I introduce a novel module specifically designed to address this issue.
+In this paper, we present an innovative solution to mitigate the entropy collapse phenomenon observed in Transformers applied to time series forecasting.
+Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts.
+Although we do not exhaustively explore all potential causes of entropy collapse, we introduce a combination of novel modules specifically combined to address this issue.
 Looking ahead, future research will focus on a comprehensive investigation of the remaining factors contributing to entropy collapse.
 This exploration aims to identify additional modifications and enhancements that could further improve the stability and performance of Transformer models in time series forecasting.
-In summary, I believe that our work not only addresses critical challenges associated with the entropy collapse phenomenon but also sets the stage for future advancements in time series forecasting methodologies.
+In summary, we believe that our work not only addresses critical challenges associated with the entropy collapse phenomenon but also sets the stage for future advancements in time series forecasting methodologies.

### v4

@@
-LSEAttention is All You Need for Time Series Forecasting
+LATST: Are Transformers Necessarily Complex for Time-Series Forecasting
 Abstract
@@
 We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models.
-Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
+Notably, LATST achieves competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
 1 Introduction
@@
 Recent research has increasingly focused on applying Transformer architectures to time-series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms.
-However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Zenget al.
+However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Ni et al.
 [1], raising concerns about the efficacy of Transformers in this context, particularly given their suboptimal performance and inherent complexity.
 1.1 Transformer-based Long-term Time Series Forecasting
-The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model which established its potential as a universal backbone for various tasks across multiple modalities.[2] The attention mechanism at the core of this architecture is formulated as follows:
+The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model [3] which established its potential as a universal backbone for various tasks across multiple modalities.
+The attention mechanism at the core of this architecture is formulated as follows:
 $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
@@
 Recent studies have tailored Transformer models specifically for long-term time-series forecasting.
-For example, Informer introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.[3] Similarly, Autoformer employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.[4]
-FEDformer incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.[5] Instead of concentrating solely on point-wise attention, PatchTST emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.[6]
-The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert in their SAMformer model.
-SAMformer integrates Reversible Instance Normalization (RevIN) to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.et al.
-[8, 7]This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
+For example, Informer [4] introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.
+Similarly, Autoformer [5] employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.
+FEDformer [6] incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.
+Instead of concentrating solely on point-wise attention, PatchTST [7] emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.
+The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert et al.
+[8] in their SAMformer model.
+SAMformer integrates Reversible Instance Normalization (RevIN) [9] to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.
+This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
 1.2 Entropy Collapse
@@
 In this work, we present the following contributions:
-  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the attention for time-series data.
-  • Introduction of the LATST (LESAttention Time-Series Transformer) a Transformer with LSEAttention replacement that attempts to solve the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
-  • Empirical validation of our approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art Transformers.
+  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the softmax operation.
+  • Introduction of the LSEAttention module, a variation of traditional multi-head attention that addresses the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
+  • Empirical validation of my approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art models.
 2 Proposed Approach
@@
 Transformers have emerged as a powerful architecture for capturing temporal associations in sequential data, primarily due to their reliance on point-wise attention mechanisms.
-However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly identical matrix.
-This identity diminishes the model’s ability to differentiate patterns, thereby hindering generalization.
+However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly uniform distributions across diverse inputs.
+This uniformity diminishes the model’s ability to differentiate patterns, thereby hindering generalization.
 An equally critical yet less-explored issue is entropy collapse, particularly prevalent in the application of Transformers to time-series forecasting.
-In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation.[1] A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
-In the attention mechanism Eq.
-1, attention scores are computed as the dot product of query and key vectors.
+In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation [1].
+A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
+In the attention mechanism (Equation 1), attention scores are computed as the dot product of query and key vectors.
 This operation tends to disproportionately emphasize abrupt changes or noise by assigning them larger values.
@@
 the exponential computation introduces significant numerical instability.
-Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.[10] This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation.
-Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output.
+Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.
+This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation.
+Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output [11].
 This disproportionate weighting undermines the diversity of the attention mechanism, reducing its capacity to effectively model complex temporal relationships.
-As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data which is the inherent characteristic of time-series data.
+As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data.
 This behavior culminates in entropy collapse, where the attention distribution becomes overly concentrated on a few elements, thereby limiting the model’s representational power.
-In other words, trend and periodicity might be barely considered.[1] Addressing this issue necessitates exploring alternatives to the conventional softmax function.
+In other words, trend and periodicity are barely considered.
+Addressing this issue necessitates exploring alternatives to the conventional softmax function.
 $\text{LSE}(x)=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right).$ (4)
-To address these challenges, we propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eq.
-4 trick proposed by Blanchard et al.
+To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eqn4 trick proposed by Blanchard et al.
 [11] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al.
-[11] to form a gated attention .
+[12].
 The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value.
-Starting from LSE where the softmax can be represented with LSE in this way:
+By starting from LSE where the softmax can be represented with LSE in this way:
 $\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
-where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing the numerical stability.
+where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing numerical stability.
 The LSE trick is stated as:
-$y=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)=\log\left(a\sum_{i=1}^{n}e^{x_{i}}-a\right)$ (6)
-where $a$ is a constant and maximum value utilized ensuring that the largest exponent becomes zero and the others become non-positive.
-As a result, all exponential terms are bounded between 0 and 1 like normalization for numerical stabilization purpose in practice.
+$y=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)=\log\left(e^{a}\sum_{i=1}^{n}e^{x_{i}-a}\right)$ (6)
+where $a$ is a constant and max value utilized for ensuring that the largest exponent becomes zero and the others become non-positive.
+As a result, all exponential terms are bounded between 0 and 1 like normalization in practice.
 This can be simplified to:
 $y=a+\log\left(\sum_{i=1}^{n}e^{x_{i}-a}\right)$ (7)
-To enhance the model’s capabitlity, the GELU is applied to introduce the non-linearity, then the equation is modfied as:
+To enhance the model’s capabitlity, the GELU is applied to indroduce the non-linearity, then the equation is modfied as:
 $y=\text{GELU}\left(a+\log\left(\sum_{i=1}^{n}e^{x_{i}-a}\right)\right)$ (8)
 This eventually can be expressed as the numerically stable manner of softmax function:
-$g_{i}=\exp(x_{i}-y)$ (9)
+$g_{i}=\exp(x_{i}-y)=\frac{e^{x_{i}}}{e^{y}}$ (9)
+$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=\text{softmax}(x_{i})$ (10)
 where $g_{i}$ represents the $i$-th component of the stabilized softmax output.
@@
 This normalization technique is particularly effective for mitigating discrepancies between training and testing data distributions in time series forecasting tasks as in [8]
-The architecture replaces the traditional temporal self-attention mechanism, with the LSEAttention module.
+The architecture retains a traditional temporal self-attention mechanism, integral to the LSEAttention module.
 To assess the efficacy of the LSEAttention component, We conducted an ablation study comparing LATST with the Gaussian Error linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified linear Unit (ReLU).
@@
 3.
-Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing adaptive learning while maintaining the robustness of attention mechanisms.
+Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing for adaptive learning while maintaining the robustness of attention mechanisms.
 This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
-Figure 1: Overall structure of LATST.
+Figure 1: Overall structure of LATST
 3 Experiments
@@
 Evaluation Metric: Consistent with previous studies, We employ the widely used evaluation metric, mean squared error (MSE).
-This metric has been a standard choice in earlier works and is defined as follows:
-$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(Y_{i}-\hat{Y}_{i})^{2},$ (10)
 where $N$ is the total number of samples, $Y_{i}$ is the $i$-th prediction, and $\hat{Y}_{i}$ represents the corresponding ground truth value.
@@
 Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
-Table 1: Hyperparameter settings used in experiments.
 Hyperparameter Values
 Batch size 8
-Initial learning rate 0.5 to 0.00005
+Initial learning rate 0.5 to 0.00001
 Number of heads 4 to 16
@@
 Prediction length 96, 192, 336, 720
+Table 1: Hyperparameter settings used in experiments.
 Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
@@
 Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
-ECL 96 0.131 0.155 0.182 0.173 - 0.140
+Electricity 96 0.131 0.155 0.182 0.173 - 0.140
 192 0.154 0.168 0.202 0.204 - 0.153
@@
 720 0.865 1.003 1.323 1.204 0.847 0.603
-Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
+Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
+Electricity Avg 0.163 0.181 0.209 0.209 - 0.166
+Traffic Avg 0.415 0.425 0.457 0.620 0.428 0.434
+Weather Avg 0.233 0.260 0.279 0.267 0.258 0.246
+ETTh1 Avg 0.414 0.410 0.554 0.439 0.454 0.423
+ETTh2 Avg 0.355 0.344 0.412 0.357 0.383 0.431
+ETTm1 Avg 0.363 0.373 0.427 0.385 0.407 0.357
+ETTm2 Avg 0.266 0.269 0.302 0.289 0.288 0.267
+Exchange Avg 0.412 0.445 0.620 0.593 0.360 0.287
 Methods LATST SAMformer
 Prediction Length 96 192 336 720 96 192 336 720
-Electricity 251K 263K 282K 331K 82K 131K 205K 401K
-Traffic 251K 263K 282K 331K 82K 131K 205K 401K
-Weather 51K 54K 58K 72K 82K 131K 205K 401K
-ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
-ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
-ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
-ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
-Exchange 51K 115K 199K 236K 82K 131K 205K 401K
+Electricity Avg 251K 263K 282K 331K 82K 131K 205K 401K
+Traffic Avg 251K 263K 282K 331K 82K 131K 205K 401K
+Weather Avg 50.5K 53.7K 58.5K 71.7K 82K 131K 205K 401K
+ETT Avg - - - - 82K 131K 205K 401K
+Exchange Avg - 115K - 236K 82K 131K 205K 401K
 3.3 Performance Analysis
-LATST demonstrates exceptional performance, achieving the lowest mean squared error (MSE) scores in 21 out of 32 scenarios across eight diverse datasets.
-This remarkable consistency underscores the model’s robustness and adaptability, effectively addressing a wide range of time-series forecasting tasks.
-Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, demonstrating its superior ability to manage complex temporal dependencies.
-When comparing LATST directly with SAMformer, the model shows its superiority by outperforming SAMformer on 6 out of 8 datasets with overall improvement by 5.12% .
-Moreover, LATST surpasses the traditional Transformer model by a notable margin of 19.75%, which emphasizes its enhanced predictive capabilities.
-Key innovations in LATST, such as the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute significantly to this performance enhancement.
-These modifications enable LATST to capture long-range dependencies and temporal patterns more effectively, setting it apart from its counterparts.
+LATST demonstrates superior performance by achieving the best mean squared error (MSE) scores in 21 out of 32 scenarios across the evaluated eight datasets.
+This remarkable consistency underscores the model’s robustness and adaptability in addressing diverse time series forecasting tasks.
+Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, indicating its efficacy in managing complex temporal dependencies.
+When comparing LATST directly with SAMformer, LATST exhibits an overall improvement of 4.3% in MSE.
+Furthermore, LATST surpasses the Transformer model by a substantial margin of 19.11%, highlighting its enhanced predictive capabilities.
+The modifications introduced in LATST, including the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute to this performance enhancement, allowing the model to better capture long-range dependencies and temporal patterns.
 The detailed performance metrics for each dataset and prediction length are summarized in Tab.
 2.
-Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variants.
-This phenomenon is closely related to the number of variants and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions.
-A more in-depth analysis can be found in Tab.
-3.
-In comparison with DLinear, LATST outperforms the model on 5 out of 8 datasets, despite DLinear having 139.7K parameters.
-LATST, with fewer parameters in most scenarios, demonstrates that its performance is not solely dependent on model size, but rather on the efficiency and effectiveness of its architecture.
-Consistency Across Datasets: LATST’s ability to achieve the best MSE scores in a significant number of scenarios further reinforces its robustness and generalizability across diverse forecasting challenges.
-This performance stability across different datasets suggests that LATST effectively learns and captures the underlying temporal patterns, allowing it to adapt well to varying data distributions and dynamics.
-For instance, in the Traffic dataset, LATST consistently maintains lower MSE values even at longer prediction horizons, illustrating its resilience against overfitting and its capacity to generalize beyond training conditions.
-Table 4: Performance of LATST with different activation functions for prediction length of 96.
-Dataset With PReLU With ReLU
-ETTh2 0.290 0.292
-ETTm1 0.295 0.305
-ETTm2 0.176 0.191
-Exchange Rate 0.098 0.115
+Notably, LATST consistently achieves the lowest MSE across multiple input lengths for datasets such as ETTh1, ETTh2, and Electricity.
+For instance, in the Electricity dataset, LATST records an MSE of 0.131 for a prediction length of 96, significantly outperforming the closest competitor, SAMformer, which achieves an MSE of 0.155.
+This trend is replicated across various prediction horizons, demonstrating the model’s ability to maintain stability and robustness under different forecasting conditions.
+Consistency Across Datasets: The model’s ability to achieve the best MSE scores in a significant number of scenarios reinforces its robustness and generalizability to diverse forecasting challenges.
+The performance stability observed across different datasets suggests that LATST effectively learns and captures underlying temporal patterns, enabling it to adapt well to varying data distributions and dynamics.
+For example, in the Traffic dataset, LATST maintains lower MSE values even at longer prediction horizons, indicating its resilience against overfitting and its capacity to generalize beyond training conditions.
 3.4 Ablation Study
 In this section, We present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework.
-By systematically removing or modifying key elements, such as the GeLU and the PReLU, their contributions are isolated to the overall performance of the model.
+By systematically removing or modifying key elements, such as the GeLU and the PReLU, We isolate their contributions to the overall performance of the model.
 For consistency, We fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
-Table 5: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
-Dataset Without GeLU Transformer
-ETTh2 0.312 0.396
-ETTm1 0.333 0.384
-ETTm2 0.183 0.200
-Exchange Rate 0.105 0.292
-Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
+Dataset Pred. Length With PReLU With ReLU
+ETTh2 96 0.290 0.292
+ETTm1 96 0.300 0.305
+ETTm2 96 0.176 0.191
+Exchange Rate 96 0.098 0.115
+Table 3: Performance of LATST with different activation functions
+Preliminary results from the ablation study, as shown in Tab.
+3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
 This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
 To further investigate the contribution of the Log-Sum-Exp trick, We conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
-As illustrated in LABEL:tab:ab_gelu, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
+Dataset Pred. Length Without GeLU Transformer
+ETTh2 96 0.312 0.396
+ETTm1 96 0.333 0.384
+ETTm2 96 0.183 0.200
+Exchange Rate 96 0.105 0.292
+Table 4: Performance of LATST without GeLU compared to Transformer
+As illustrated in Tab.
+4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
 This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function.
@@
 Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts.
-Although we do not exhaustively explore all potential causes of entropy collapse, we introduce a combination of novel modules specifically combined to address this issue.
+Although we do not exhaustively explore all potential causes of entropy collapse, We introduce a novel module specifically designed to address this issue.
 Looking ahead, future research will focus on a comprehensive investigation of the remaining factors contributing to entropy collapse.

### v5

@@
-LATST: Are Transformers Necessarily Complex for Time-Series Forecasting
+LSEAttention is All You Need for Time Series Forecasting
 Abstract
@@
 We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models.
-Notably, LATST achieves competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
+Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
 1 Introduction
@@
 Recent research has increasingly focused on applying Transformer architectures to time-series forecasting, leveraging their capacity to model complex temporal interactions through self-attention mechanisms.
-However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Ni et al.
+However, despite their potential, state-of-the-art approaches to multivariate time series forecasting still rely mainly on linear models, such as those proposed by Zenget al.
 [1], raising concerns about the efficacy of Transformers in this context, particularly given their suboptimal performance and inherent complexity.
 1.1 Transformer-based Long-term Time Series Forecasting
-The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model [3] which established its potential as a universal backbone for various tasks across multiple modalities.
-The attention mechanism at the core of this architecture is formulated as follows:
+The application of Transformers to time series forecasting gained momentum following the introduction of the Transformer model which established its potential as a universal backbone for various tasks across multiple modalities.[2] The attention mechanism at the core of this architecture is formulated as follows:
 $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (1)
@@
 Recent studies have tailored Transformer models specifically for long-term time-series forecasting.
-For example, Informer [4] introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.
-Similarly, Autoformer [5] employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.
-FEDformer [6] incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.
-Instead of concentrating solely on point-wise attention, PatchTST [7] emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.
-The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert et al.
-[8] in their SAMformer model.
-SAMformer integrates Reversible Instance Normalization (RevIN) [9] to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.
-This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
+For example, Informer introduces ProbSparse self-attention and distillation techniques to efficiently identify the most relevant keys to forecasting, thus reducing computational complexity and improving interpretability by focusing on essential temporal signals.[3] Similarly, Autoformer employs decomposition and auto-correlation mechanisms, effectively merging classical and modern methodologies to enhance model performance.[4]
+FEDformer incorporates a Fourier-enhanced module that achieves linear complexity in both time and space, significantly improving scalability and efficiency for long sequences.[5] Instead of concentrating solely on point-wise attention, PatchTST emphasizes patch-level attention by treating patches, rather than individual time steps, as input units, allowing the model to capture richer semantic information across multiple time series - a critical aspect for effective long-term forecasting.[6]
+The challenges associated with applying Transformers to time series forecasting are further highlighted by Ilbert in their SAMformer model.
+SAMformer integrates Reversible Instance Normalization (RevIN) to address shifting data distributions and incorporates Sharpness-Aware Minimization (SAM) to improve training stability.et al.
+[8, 7]This approach tackles critical issues, including entropy collapse and training instability, that have impeded Transformer models in this domain.
 1.2 Entropy Collapse
@@
 In this work, we present the following contributions:
-  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the softmax operation.
-  • Introduction of the LSEAttention module, a variation of traditional multi-head attention that addresses the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
-  • Empirical validation of my approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art models.
+  • A novel theoretical framework linking entropy collapse with numerical stability, aimed at restoring the intended functionality of the attention for time-series data.
+  • Introduction of the LATST (LESAttention Time-Series Transformer) a Transformer with LSEAttention replacement that attempts to solve the entropy collapse issue, thereby enhancing the trainability and performance of Transformers for time series forecasting.
+  • Empirical validation of our approach on widely used multivariate long-term forecasting datasets, demonstrating that Transformers equipped with LSEAttention outperform state-of-the-art Transformers.
 2 Proposed Approach
@@
 Transformers have emerged as a powerful architecture for capturing temporal associations in sequential data, primarily due to their reliance on point-wise attention mechanisms.
-However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly uniform distributions across diverse inputs.
-This uniformity diminishes the model’s ability to differentiate patterns, thereby hindering generalization.
+However, this reliance introduces vulnerabilities, including a phenomenon referred to as attention collapse, where attention matrices converge to nearly identical matrix.
+This identity diminishes the model’s ability to differentiate patterns, thereby hindering generalization.
 An equally critical yet less-explored issue is entropy collapse, particularly prevalent in the application of Transformers to time-series forecasting.
-In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation [1].
-A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
-In the attention mechanism (Equation 1), attention scores are computed as the dot product of query and key vectors.
+In such scenarios, Transformers might overfit to transient noise or abrupt changes in the data, resulting in significant performance degradation.[1] A deeper investigation into these phenomena reveals that the root cause often lies in the computation of the softmax function, especially the handling of its exponential terms.
+In the attention mechanism Eq.
+1, attention scores are computed as the dot product of query and key vectors.
 This operation tends to disproportionately emphasize abrupt changes or noise by assigning them larger values.
@@
 The exponential computation introduces significant numerical instability.
-Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.
-This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation.
-Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output [11].
+Specifically, large attention scores lead to exponential overflow, while small attention scores result in underflow.[10] This imbalance is further exacerbated during normalization, as the largest attention score (corresponding to the overflowed exponential value) disproportionately dominates the computation.
+Conversely, smaller attention scores, which result from underflowed exponential values, are effectively suppressed and contribute negligibly to the output.
 This disproportionate weighting undermines the diversity of the attention mechanism, reducing its capacity to effectively model complex temporal relationships.
-As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data.
+As a consequence, the model’s ability to generalize is compromised, and its performance deteriorates in the presence of noisy or volatile data which is the inherent characteristic of time-series data.
 This behavior culminates in entropy collapse, where the attention distribution becomes overly concentrated on a few elements, thereby limiting the model’s representational power.
-In other words, trend and periodicity are barely considered.
-Addressing this issue necessitates exploring alternatives to the conventional softmax function.
+In other words, trend and periodicity might be barely considered.[1] Addressing this issue necessitates exploring alternatives to the conventional softmax function.
 $\text{LSE}(x)=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right).$ (4)
@@
 This normalization technique is particularly effective for mitigating discrepancies between training and testing data distributions in time series forecasting tasks as in [7]
-The architecture retains a traditional temporal self-attention mechanism, integral to the LSEAttention module.
+The architecture replaces the traditional temporal self-attention mechanism, with the LSEAttention module.
 To assess the efficacy of the LSEAttention component, we conducted an ablation study comparing LATST with the Gaussian Error linear Unit (GeLU) in the Feedforward Neural Network (FFN) against a variant utilizing the Rectified linear Unit (ReLU).
@@
 3.
-Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing for adaptive learning while maintaining the robustness of attention mechanisms.
+Overall, the LATST architecture consists of a single-layer Transformer framework augmented with substitution modules, allowing adaptive learning while maintaining the robustness of attention mechanisms.
 This design facilitates effective modeling of temporal dependencies and enhances performance in time series forecasting tasks.
-Figure 1: Overall structure of LATST
+Figure 1: Overall structure of LATST.
 3 Experiments
@@
 This flexibility in prediction length allows for an evaluation of the model’s performance across different forecasting horizons.
-Evaluation Metric: Consistent with previous studies, We employ the widely used evaluation metric, mean squared error (MSE).
-where $N$ is the total number of samples, $Y_{i}$ is the $i$-th prediction, and $\hat{Y}_{i}$ represents the corresponding ground truth value.
+Evaluation Metric: Consistent with previous studies, we employ the widely used evaluation metric, mean squared error (MSE) and mean absolute error (MAE).
+This metric has been a standard choice in earlier works.
 3.2 Baseline
@@
 Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
+Table 1: Hyperparameter settings used in experiments.
 Hyperparameter Values
 Batch size 8
-Initial learning rate 0.5 to 0.00001
+Initial learning rate 0.5 to 0.00005
 Number of heads 4 to 16
@@
 Prediction length 96, 192, 336, 720
-Table 1: Hyperparameter settings used in experiments.
 Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
 The best results are highlighted in bold.
-LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while results for other models are sourced from Ilbert et al.
+LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while most results for other models are sourced from Ilbert et al.
+[7] with some ECL result from Liu et al.
 [16].
 Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
-Electricity 96 0.131 0.155 0.182 0.173 - 0.140
-192 0.154 0.168 0.202 0.204 - 0.153
-336 0.169 0.183 0.212 0.217 - 0.169
-720 0.197 0.219 0.238 0.242 - 0.203
+ECL 96 0.131 0.155 0.182 0.173 0.148 0.140
+192 0.154 0.168 0.202 0.204 0.162 0.153
+336 0.169 0.183 0.212 0.217 0.178 0.169
+720 0.197 0.219 0.238 0.242 0.225 0.203
 Traffic 96 0.380 0.407 0.420 0.409 0.395 0.410
@@
 720 0.865 1.003 1.323 1.204 0.847 0.603
-Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
-Electricity Avg 0.163 0.181 0.209 0.209 - 0.166
-Traffic Avg 0.415 0.425 0.457 0.620 0.428 0.434
-Weather Avg 0.233 0.260 0.279 0.267 0.258 0.246
-ETTh1 Avg 0.414 0.410 0.554 0.439 0.454 0.423
-ETTh2 Avg 0.355 0.344 0.412 0.357 0.383 0.431
-ETTm1 Avg 0.363 0.373 0.427 0.385 0.407 0.357
-ETTm2 Avg 0.266 0.269 0.302 0.289 0.288 0.267
-Exchange Avg 0.412 0.445 0.620 0.593 0.360 0.287
+Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
 Methods LATST SAMformer
 Prediction Length 96 192 336 720 96 192 336 720
-Electricity Avg 251K 263K 282K 331K 82K 131K 205K 401K
-Traffic Avg 251K 263K 282K 331K 82K 131K 205K 401K
-Weather Avg 50.5K 53.7K 58.5K 71.7K 82K 131K 205K 401K
-ETT Avg - - - - 82K 131K 205K 401K
-Exchange Avg - 115K - 236K 82K 131K 205K 401K
+Electricity 251K 263K 282K 331K 82K 131K 205K 401K
+Traffic 251K 263K 282K 331K 82K 131K 205K 401K
+Weather 51K 54K 58K 72K 82K 131K 205K 401K
+ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
+ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
+ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
+ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
+Exchange 51K 115K 199K 236K 82K 131K 205K 401K
 3.3 Performance Analysis
-LATST demonstrates superior performance by achieving the best mean squared error (MSE) scores in 21 out of 32 scenarios across the evaluated eight datasets.
-This remarkable consistency underscores the model’s robustness and adaptability in addressing diverse time series forecasting tasks.
-Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, indicating its efficacy in managing complex temporal dependencies.
-When comparing LATST directly with SAMformer, LATST exhibits an overall improvement of 4.3% in MSE.
-Furthermore, LATST surpasses the Transformer model by a substantial margin of 19.11%, highlighting its enhanced predictive capabilities.
-The modifications introduced in LATST, including the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute to this performance enhancement, allowing the model to better capture long-range dependencies and temporal patterns.
+LATST demonstrates exceptional performance, achieving the lowest mean squared error (MSE) scores in 21 out of 32 scenarios across eight diverse datasets.
+This remarkable consistency underscores the model’s robustness and adaptability, effectively addressing a wide range of time-series forecasting tasks.
+Specifically, LATST outperforms all other Transformer-based models on the Electricity, Traffic, and Weather datasets across all evaluated metrics, demonstrating its superior ability to manage complex temporal dependencies.
+When comparing LATST directly with SAMformer, the model shows its superiority by outperforming SAMformer on 6 out of 8 datasets with overall improvement by 4% .
+Moreover, LATST surpasses the traditional Transformer model by a notable margin of 19.75%, which emphasizes its enhanced predictive capabilities.
+Key innovations in LATST, such as the integration of the Log-Sum-Exp trick and refined attention mechanisms, contribute significantly to this performance enhancement.
+These modifications enable LATST to capture long-range dependencies and temporal patterns more effectively, setting it apart from its counterparts.
 The detailed performance metrics for each dataset and prediction length are summarized in Tab.
 2.
-Notably, LATST consistently achieves the lowest MSE across multiple input lengths for datasets such as ETTh1, ETTh2, and Electricity.
-For instance, in the Electricity dataset, LATST records an MSE of 0.131 for a prediction length of 96, significantly outperforming the closest competitor, SAMformer, which achieves an MSE of 0.155.
-This trend is replicated across various prediction horizons, demonstrating the model’s ability to maintain stability and robustness under different forecasting conditions.
-Consistency Across Datasets: The model’s ability to achieve the best MSE scores in a significant number of scenarios reinforces its robustness and generalizability to diverse forecasting challenges.
-The performance stability observed across different datasets suggests that LATST effectively learns and captures underlying temporal patterns, enabling it to adapt well to varying data distributions and dynamics.
-For example, in the Traffic dataset, LATST maintains lower MSE values even at longer prediction horizons, indicating its resilience against overfitting and its capacity to generalize beyond training conditions.
+Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variables.
+This phenomenon is closely related to the number of variables and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions.
+A more in-depth analysis can be found in Tab.
+3.
+In comparison with DLinear, LATST outperforms the model on 5 out of 8 datasets, despite DLinear having 139.7K parameters.
+LATST, with fewer parameters in most scenarios, demonstrates that its performance is not solely dependent on model size, but rather on the efficiency and effectiveness of its architecture.
+Consistency Across Datasets: LATST’s ability to achieve the best MSE scores in a significant number of scenarios further reinforces its robustness and generalizability across diverse forecasting challenges.
+This performance stability across different datasets suggests that LATST effectively learns and captures the underlying temporal patterns, allowing it to adapt well to varying data distributions and dynamics.
+For instance, in the Traffic dataset, LATST consistently maintains lower MSE values even at longer prediction horizons, illustrating its resilience against overfitting and its capacity to generalize beyond training conditions.
+Table 4: Performance of LATST with different activation functions for prediction length of 96.
+Dataset With PReLU With ReLU
+ETTh2 0.290 0.292
+ETTm1 0.295 0.305
+ETTm2 0.176 0.191
+Exchange Rate 0.098 0.115
 3.4 Ablation Study
 In this section, we present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework.
-By systematically removing or modifying key elements, such as the GeLU and the PReLU, We isolate their contributions to the overall performance of the model.
+By systematically removing or modifying key elements, such as the GeLU and the PReLU, their contributions are isolated to the overall performance of the model.
 For consistency, we fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
-Dataset Pred. Length With PReLU With ReLU
-ETTh2 96 0.290 0.292
-ETTm1 96 0.300 0.305
-ETTm2 96 0.176 0.191
-Exchange Rate 96 0.098 0.115
-Table 3: Performance of LATST with different activation functions
-Preliminary results from the ablation study, as shown in Tab.
-3, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
+Table 5: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
+Dataset Without GeLU Transformer
+ETTh2 0.312 0.396
+ETTm1 0.333 0.384
+ETTm2 0.183 0.200
+Exchange Rate 0.105 0.292
+Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.
 This observation suggests that the performance gains associated with PReLU may not arise solely from its inherent properties; rather, it appears that the ReLU activation function may hinder the performance of the LSEAttention model due to the numerical instability introduced by its sharp activation characteristics.
 To further investigate the contribution of the Log-Sum-Exp trick, we conducted an experiment comparing LATST with a Transformer model that replaces the GeLU activation function.
-Dataset Pred. Length Without GeLU Transformer
-ETTh2 96 0.312 0.396
-ETTm1 96 0.333 0.384
-ETTm2 96 0.183 0.200
-Exchange Rate 96 0.105 0.292
-Table 4: Performance of LATST without GeLU compared to Transformer
-As illustrated in Tab.
-4, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
+As illustrated in LABEL:tab:ab_gelu, the LATST model utilizing the Log-Sum-Exp trick significantly outperforms the Transformer model, despite a notable decline in performance compared to the original configuration.
 This indicates that the Log-Sum-Exp trick plays a crucial role in maintaining the effectiveness of the LATST architecture, even in the absence of the GeLU activation function.
@@
 Our empirical results confirm the effectiveness and robustness of the proposed LATST approach inty of forecasting contexts.
-Although we do not exhaustively explore all potential causes of entropy collapse, We introduce a novel module specifically designed to address this issue.
+Although we do not exhaustively explore all potential causes of entropy collapse, we introduce a combination of novel modules specifically combined to address this issue.
 Looking ahead, future research will focus on a comprehensive investigation of the remaining factors contributing to entropy collapse.

### v6

@@
 $\text{LSE}(x)=\log\left(\sum_{i=1}^{n}e^{x_{i}}\right).$ (4)
-To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eqn4 trick proposed by Blanchard et al.
+To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eq.
+4 trick proposed by Blanchard et al.
 [10] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al.
@@
 By starting from LSE where the softmax can be represented with LSE in this way:
-$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
+$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{i=1}^{n}e^{x_{i}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
 where $e^{\text{LSE}(x)}$ denotes the exponential of the log-sum-exp function, enhancing numerical stability.
@@
 $g_{i}=\exp(x_{i}-y)=\frac{e^{x_{i}}}{e^{y}}$ (9)
-$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=\text{softmax}(x_{i})$ (10)
+$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{i=1}^{n}e^{x_{i}}}=\text{softmax}(x_{i})$ (10)
 where $g_{i}$ represents the $i$-th component of the stabilized softmax output.
+Instead of strictly converting back to original softmax function in Eq.
+10.
+We instead of applying softmax to the $g_{i}$ to normalize the final attention weights into the bound between 0 and 1 in Eq.
+11.
+$\text{softmax}(g_{i})$ (11)
 In traditional Transformer architectures, the ReLU (Rectified linear Unit) activation function used in the Feed-Forward Network (FFN) is susceptible to the ”dying ReLU” problem, where neurons can become inactive by outputting zero for all negative input values.
@@
 Each time series is segmented based on a specified input length to facilitate effective modeling.
-Sequence and Prediction Length: In accordance with prior works, We set the input sequence length at 336 time steps.
+Sequence and Prediction Length: In accordance with prior works, We set the input sequence length at 512 time steps for all models.
 The prediction length, which indicates the number of future time steps for which We aim to forecast data, varies among the set {96, 192, 336, 720}.
@@
 Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
-Table 1: Hyperparameter settings used in experiments.
-Hyperparameter Values
-Batch size 8
-Initial learning rate 0.5 to 0.00005
-Number of heads 4 to 16
-Head dropout rate 0, 0.2, 0.3
-Input sequence length 336
-Prediction length 96, 192, 336, 720
+Rest of the hyparameters could be referred in Tab.
+2 where it demonsteates clear pattern between the set of parameters and the categories.
 Table 1: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
@@
 [16].
-Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
-ECL 96 0.131 0.155 0.182 0.173 0.148 0.140
-192 0.154 0.168 0.202 0.204 0.162 0.153
-336 0.169 0.183 0.212 0.217 0.178 0.169
-720 0.197 0.219 0.238 0.242 0.225 0.203
-Traffic 96 0.380 0.407 0.420 0.409 0.395 0.410
-192 0.406 0.415 0.441 0.637 0.417 0.423
-336 0.417 0.421 0.501 0.747 0.433 0.436
-720 0.455 0.456 0.468 0.688 0.467 0.466
-Weather 96 0.152 0.197 0.227 0.214 0.174 0.176
-192 0.196 0.235 0.256 0.231 0.221 0.220
-336 0.249 0.276 0.278 0.279 0.278 0.265
-720 0.317 0.334 0.353 0.343 0.358 0.323
-ETTh1 96 0.390 0.381 0.509 0.398 0.386 0.375
-192 0.426 0.409 0.535 0.426 0.441 0.405
-336 0.414 0.423 0.570 0.435 0.487 0.439
-720 0.426 0.427 0.601 0.498 0.503 0.472
-ETTh2 96 0.290 0.295 0.396 0.308 0.297 0.289
-192 0.348 0.340 0.413 0.352 0.380 0.383
-336 0.380 0.350 0.414 0.360 0.428 0.448
-720 0.399 0.391 0.424 0.409 0.427 0.605
-ETTm1 96 0.295 0.329 0.384 0.336 0.334 0.299
-192 0.345 0.353 0.400 0.362 0.377 0.335
-336 0.375 0.382 0.461 0.391 0.426 0.369
-720 0.431 0.429 0.463 0.450 0.491 0.425
-ETTm2 96 0.176 0.181 0.200 0.211 0.180 0.167
-192 0.229 0.233 0.273 0.252 0.250 0.224
-336 0.280 0.285 0.310 0.303 0.311 0.281
-720 0.376 0.375 0.426 0.390 0.412 0.397
-Exchange 96 0.098 0.161 0.292 0.343 0.086 0.081
-192 0.218 0.246 0.372 0.342 0.177 0.157
-336 0.408 0.368 0.494 0.484 0.331 0.305
-720 0.865 1.003 1.323 1.204 0.847 0.603
-Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
-Methods LATST SAMformer
-Prediction Length 96 192 336 720 96 192 336 720
-Electricity 251K 263K 282K 331K 82K 131K 205K 401K
-Traffic 251K 263K 282K 331K 82K 131K 205K 401K
-Weather 51K 54K 58K 72K 82K 131K 205K 401K
-ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
-ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
-ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
-ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
-Exchange 51K 115K 199K 236K 82K 131K 205K 401K
+Methods LATST SAMformer PatchTST Transformer TSMixer iTrans Dlinear
+ECL 96 0.130 0.155 - 0.182 0.173 - 0.140
+192 0.151 0.168 - 0.202 0.204 - 0.153
+336 0.169 0.183 - 0.212 0.217 - 0.169
+720 0.197 0.219 - 0.238 0.242 - 0.203
+Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
+192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
+336 0.417 0.421 0.482 0.501 0.747 0.433 0.436
+720 0.455 0.456 0.514 0.468 0.688 0.467 0.466
+Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
+192 0.196 0.235 0.225 0.256 0.231 0.221 0.220
+336 0.249 0.276 0.278 0.278 0.279 0.278 0.265
+720 0.317 0.334 0.354 0.353 0.343 0.358 0.323
+ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
+192 0.411 0.409 0.460 0.535 0.426 0.441 0.405
+336 0.433 0.423 0.501 0.570 0.435 0.487 0.439
+720 0.426 0.427 0.500 0.601 0.498 0.503 0.472
+ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
+192 0.343 0.340 0.388 0.413 0.352 0.380 0.383
+336 0.380 0.350 0.426 0.414 0.360 0.428 0.448
+720 0.399 0.391 0.431 0.424 0.409 0.427 0.605
+ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
+192 0.342 0.353 0.367 0.400 0.362 0.377 0.335
+336 0.375 0.382 0.399 0.461 0.391 0.426 0.369
+720 0.431 0.429 0.454 0.463 0.450 0.491 0.425
+ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
+192 0.230 0.233 0.241 0.273 0.252 0.250 0.224
+336 0.280 0.285 0.305 0.310 0.303 0.311 0.281
+720 0.376 0.375 0.402 0.426 0.390 0.412 0.397
+Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
+192 0.211 0.246 0.176 0.372 0.342 0.177 0.157
+336 0.408 0.368 0.301 0.494 0.484 0.331 0.305
+720 0.865 1.003 0.901 1.323 1.204 0.847 0.603
+Table 2: Classifications of Dataset where Low Variates: N = 4, D = 32, F = 64; Mid Variates: N = 8 D = 32, F = 128; High Variates: N = 16, D = 128, F = 256; Ultra-High Variates: N = 16 D = 256, F = 512.
+It demonstrates a clear pattern between number of variates and number of heads, as well as the dimension sizes.
+Dataset Type
+Low Variates ETTH1
+ETTH2
+ETTM1
+ETTM2
+Exchange
+Mid Variates Weather
+High Variates ECL
+Ultra-High Variates Traffic
 3.3 Performance Analysis
@@
 The detailed performance metrics for each dataset and prediction length are summarized in Tab.
-2.
+1.
 Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variables.
 This phenomenon is closely related to the number of variables and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions.
-A more in-depth analysis can be found in Tab.
-3.
+A more in-depth analysis can be found in LABEL:tab:parms.
 In comparison with DLinear, LATST outperforms the model on 5 out of 8 datasets, despite DLinear having 139.7K parameters.

### v8

@@
 However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines.
-Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting.
-We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models.
-Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
+Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.
+To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting.
+We rigorously evaluated LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models.
+Notably, LATST, a single-layer Transformer, manages to achieve competitive performance with fewer parameters than some linear models on average, highlighting its efficiency and effectiveness.
 1 Introduction
@@
 We compare the performance of LATST against leading Transformer-based models, including SAMFormer [7], TSMixer [15], iTransformer [16], and PatchTST [6].
-Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
+Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from SAMFomer, with a batch size fixed at 8 for simplicity.
 Rest of the hyparameters could be referred in Tab.
-2 where it demonsteates clear pattern between the set of parameters and the categories.
-Table 1: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
-The best results are highlighted in bold.
-LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while most results for other models are sourced from Ilbert et al.
-[7] with some ECL result from Liu et al.
-[16].
-Methods LATST SAMformer PatchTST Transformer TSMixer iTrans Dlinear
-ECL 96 0.130 0.155 - 0.182 0.173 - 0.140
-192 0.151 0.168 - 0.202 0.204 - 0.153
-336 0.169 0.183 - 0.212 0.217 - 0.169
-720 0.197 0.219 - 0.238 0.242 - 0.203
-Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
-192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
-336 0.417 0.421 0.482 0.501 0.747 0.433 0.436
-720 0.455 0.456 0.514 0.468 0.688 0.467 0.466
-Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
-192 0.196 0.235 0.225 0.256 0.231 0.221 0.220
-336 0.249 0.276 0.278 0.278 0.279 0.278 0.265
-720 0.317 0.334 0.354 0.353 0.343 0.358 0.323
-ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
-192 0.411 0.409 0.460 0.535 0.426 0.441 0.405
-336 0.433 0.423 0.501 0.570 0.435 0.487 0.439
-720 0.426 0.427 0.500 0.601 0.498 0.503 0.472
-ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
-192 0.343 0.340 0.388 0.413 0.352 0.380 0.383
-336 0.380 0.350 0.426 0.414 0.360 0.428 0.448
-720 0.399 0.391 0.431 0.424 0.409 0.427 0.605
-ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
-192 0.342 0.353 0.367 0.400 0.362 0.377 0.335
-336 0.375 0.382 0.399 0.461 0.391 0.426 0.369
-720 0.431 0.429 0.454 0.463 0.450 0.491 0.425
-ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
-192 0.230 0.233 0.241 0.273 0.252 0.250 0.224
-336 0.280 0.285 0.305 0.310 0.303 0.311 0.281
-720 0.376 0.375 0.402 0.426 0.390 0.412 0.397
-Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
-192 0.211 0.246 0.176 0.372 0.342 0.177 0.157
-336 0.408 0.368 0.301 0.494 0.484 0.331 0.305
-720 0.865 1.003 0.901 1.323 1.204 0.847 0.603
-Table 2: Classifications of Dataset where Low Variates: N = 4, D = 32, F = 64; Mid Variates: N = 8 D = 32, F = 128; High Variates: N = 16, D = 128, F = 256; Ultra-High Variates: N = 16 D = 256, F = 512.
-It demonstrates a clear pattern between number of variates and number of heads, as well as the dimension sizes.
-Dataset Type
+1 where it demonstrates clear pattern between the set of parameters and the categories.
+Table 1: Classifications of Datasets with corresponding combination of hyperparameters where all are single-layer with Low Variates: $N$ = 4, $D$ = 32, $F$ = 64; Mid Variates:$N$ = 8 $D$ = 32, $F$ = 128; High Variates: $N$ = 16, $D$ = 128, $F$ = 256; Ultra-High Variates: $N$ = 16 $D$ = 256, $F$ = 512.
+It demonstrates a clear positive relationship between number of variates and number of heads, as well as the dimension sizes to better capture the channel interaction.
+Dataset Type Single-Layer, LR: 0.0001
 Low Variates ETTH1
@@
 Ultra-High Variates Traffic
+Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons.
+The best results are highlighted in bold.
+LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while most results for other models are sourced from Ilbert et al.
+[7] with some ECL results from triplet layers Transformer on average from PatchTST et al.
+[6] and iTransformer et al.
+[16].
+Methods LATST SAMformer PatchTST Transformer TSMixer iTransformer Dlinear
+ECL 96 0.129 0.155 0.129 0.182 0.173 0.148 0.140
+192 0.150 0.168 0.147 0.202 0.204 0.162 0.153
+336 0.168 0.183 0.163 0.212 0.217 0.178 0.169
+720 0.205 0.219 0.197 0.238 0.242 0.225 0.203
+Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
+192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
+336 0.397 0.421 0.482 0.501 0.747 0.433 0.436
+720 0.433 0.456 0.514 0.468 0.688 0.467 0.466
+Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
+192 0.193 0.235 0.225 0.256 0.231 0.221 0.220
+336 0.246 0.276 0.278 0.278 0.279 0.278 0.265
+720 0.314 0.334 0.354 0.353 0.343 0.358 0.323
+ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
+192 0.415 0.409 0.460 0.535 0.426 0.441 0.405
+336 0.434 0.423 0.501 0.570 0.435 0.487 0.439
+720 0.471 0.427 0.500 0.601 0.498 0.503 0.472
+ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
+192 0.346 0.340 0.388 0.413 0.352 0.380 0.383
+336 0.384 0.350 0.426 0.414 0.360 0.428 0.448
+720 0.423 0.391 0.431 0.424 0.409 0.427 0.605
+ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
+192 0.341 0.353 0.367 0.400 0.362 0.377 0.335
+336 0.367 0.382 0.399 0.461 0.391 0.426 0.369
+720 0.420 0.429 0.454 0.463 0.450 0.491 0.425
+ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
+192 0.238 0.233 0.241 0.273 0.252 0.250 0.224
+336 0.287 0.285 0.305 0.310 0.303 0.311 0.281
+720 0.369 0.375 0.402 0.426 0.390 0.412 0.397
+Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
+192 0.201 0.246 0.176 0.372 0.342 0.177 0.157
+336 0.395 0.368 0.301 0.494 0.484 0.331 0.305
+720 0.935 1.003 0.901 1.323 1.204 0.847 0.603
 3.3 Performance Analysis
@@
 The detailed performance metrics for each dataset and prediction length are summarized in Tab.
-1.
+2.
 Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variables.
@@
 For instance, in the Traffic dataset, LATST consistently maintains lower MSE values even at longer prediction horizons, illustrating its resilience against overfitting and its capacity to generalize beyond training conditions.
-Table 3: Performance of LATST with different activation functions for prediction length of 96.
+Table 3: Performance of LATST with different activation functions for prediction length of 720.
 Dataset With PReLU With ReLU
-ETTh2 0.290 0.292
-ETTm1 0.295 0.305
-ETTm2 0.176 0.191
-Exchange Rate 0.098 0.115
-3.4 Ablation Study
+ECL 0.205 0.205
+Traffic 0.433 0.437
+Weather 0.314 0.314
+ETTh1 0.471 0.471
+ETTh2 0.423 0.426
+ETTm1 0.420 0.423
+ETTm2 0.369 0.369
+Exchange Rate 0.935 1.043
+3.4 Ablation Study Note Updated Yet
 In this section, we present a comprehensive ablation study aimed at quantitatively assessing the impact of various architectural components within the LATST framework.
@@
 For consistency, we fix the experimental parameters to a prediction length of 96 and utilize four datasets: ETTh2, ETTm1, ETTm2, and Exchange Rate.
-Table 4: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
-Dataset Without GeLU Transformer
-ETTh2 0.312 0.396
-ETTm1 0.333 0.384
-ETTm2 0.183 0.200
-Exchange Rate 0.105 0.292
+Table 4: Performance of LATST without GeLU compared to Transformer for prediction length of 192.
+Dataset With GeLU Without GeLU
+ECL 0.150 0.153
+Traffic 0.386 0.393
+Weather 0.193 0.204
+ETTh1 0.415 0.433
+ETTh2 0.346 0.378
+ETTm1 0.341 0.357
+ETTm2 0.238 0.246
+Exchange Rate 0.201 0.236
+Table 5: Overall performance improvement and average size of parameters with data of other models sourced from Ilbert et al. [7]
+LATST SAMformer
+MSE Error Percentage 107.17% 100.00%
+Average Size of Parameter 271.4K 173.2K
 Preliminary results from the ablation study, as shown in LABEL:tab:ab_relu, indicate that the inclusion of the PReLU activation function leads to a significant improvement on the ETTm2 dataset, while providing only marginal enhancements for the other datasets.